OpenStudy (anonymous):

10 5/6x^12+-7/24x^12

jhonyy9 (jhonyy9):

so if i see correct what you have wrote this : 5 -7 4(50) -7 10 ------ + -------- = -------- = 6x*12 24x*12 24x*12 193 = -------- 24x*12 but if you have wrote 105/6x*12 than will be : 105(4) -7 413 --------- = ------- 24x*12 24x*12